# Coherence and Transients in Nonlocally Coupled Dissipative Kicked Rotors

### v1

@@ I Introduction
However, the analysis of nonlocal coupling presents significant theoretical challenges, as it eludes the simplifying symmetries of all-to-all coupling via mean-field approaches (e.g., the Kuramoto model reduces to a single order parameter), and the spatial regularity of nearest-neighbor coupling via reaction-diffusion PDEs or discrete Laplacians. Nonlocal systems, with their distance-dependent interactions – require more sophisticated approaches. The seminal work [11] demonstrated this through coherence-incoherence transitions in nonlocally coupled logistic maps, and confirmed that such a transition is universal among systems with diverse local dynamics. Inspired by this, we present an analytical and numerical study of various transitions in a nonlocally coupled system with a two-dimensional local map, specifically the dissipative kicked rotor [12, 13, 14]. Interestingly, despite the nonlocal coupling, linear stability analysis effectively captures the dynamics of the coupled system beyond the linearization regime; the resulting phase diagram exhibits qualitatively different structure compared to previously studied systems [11]. Analyzing bifurcations of simple states reveals transition points between coherent states of different spatial periodicity, and between coherent and incoherent states. These will be illustrated in detail.
@@ II Model
where $(p_{j}(t),\theta_{j}(t))\in(-\infty,+\infty)\times[0,2\pi)$ denote the momentum and angle of the rotor $j\in\{1,2,...N\}$ at time $t\in\mathbb{N}_{0}$, $\gamma\in(0,1)$ is the dissipation coefficient, $K_{0}>0$ is the local nonlinearity parameter, $K>0$ is the coupling strength, and $P_{c}\in\left\{1,2,..,\frac{N}{2}\right\}$ (assuming $N$ is even) is the coupling length, denoting the number of neighbors in each direction that are coupled to the rotor $j$.
The local map describes a dissipative kicked rotor and for a fixed $\gamma$ the system undergoes multiple bifurcations under variations of $K_{0}$ [14]. For the nearest-neighbor case ($P_{c}=1$), see [14].
@@ III Spatial and Temporal Patterns
In this section, we focus on the following three issues. The first one is the transition from the stationary, fully synchronized state to a coherent state of temporal period-$2$, which can be explained analytically through a linear stability analysis. A wavenumber will be defined to characterize the spatial periodicity. The second issue is to study coherent regions on the parameter $(P_{c},K)$-plane, particularly the regions with low spatial and temporal periods. The transition curves can be obtained from stability analysis of the temporal period-$2$ states. The third issue concerns the coexistence of the temporal period-$2$ states with different wavenumbers, and we numerically examine how their relative basin sizes vary with the coupling length $P_{c}$.
Throughout this section, purely random initial conditions are used for numerical simulations: $(p_{j}(0),\theta_{j}(0))\in\text{Uni}[-p_{0},p_{0}]\times\text{Uni}[0,2\pi)$, where $p_{0}>2\pi$ is constant (the specific value of $p_{0}$ does not influence the qualitative results).
@@ III.1 Transition from zero homogeneous states
The full phase space of the system (1) is a product of $N$ infinite cylinders $(p_{j},\theta_{j})\in(-\infty,+\infty)\times[0,2\pi)$, $\forall j=1,2,...,N$. For $K_{0}$ and $K$ both small, trajectories initially close to $\boldsymbol{p}:=(p_{1},p_{2},...,p_{N})=\boldsymbol{0}$ will converge to the zero synchronized state $(\boldsymbol{p},\boldsymbol{\theta})=(\boldsymbol{0},\boldsymbol{0})$ and it appears as the only attractor. However, due to multistability of the local map [14], more attractors can be observed when we consider longer truncated cylinders (i.e., $\boldsymbol{p}$ further away from $\boldsymbol{0}$), and the notion of the zero synchronized state is extended to a state in which most rotors are homogeneous but with a few exceptions; these exceptional rotors stay at the nonzero fixed points of the local map. We refer such a state a zero homogeneous state. An example is shown in Fig.1(a): for $K_{0}=2$ the local map has three fixed points $p=0,\pm 2\pi$, where $p=0$ has the largest basin and ±$\pm 2\pi$ have equal basin sizes [14].
@@
Figure 1: Snapshots of momenta for (a) $K=1.3$, (b) $K=1.4$ and (c) $K=3$. Other parameters: $\gamma=0.8$, $K_{0}=2$, $P_{c}=32$, $N=100$ and random initial conditions $(p_{j}(0),\theta_{j}(0))\in\text{Uni}[-35,35]\times\text{Uni}[0,2\pi)$, $j=1,2,...,N$. The corresponding angles behave similarly.
This period-doubling bifurcation can be understood analytically via a linear stability analysis. Under the assumption of slow angle variations, the equations of momenta can be linearized as
By a Fourier transform $p_{j}(t)=\sum_{w}P_{w}(t)e^{iwj}$, $\theta_{j}(t)=\sum_{w}\Theta_{w}(t)e^{iwj}$, $w=\frac{2\pi l}{N}$, $l=0,1,...,N-1$ (for periodic boundary conditions) we have, for each Fourier mode $w$,
The linearized equations of motion thus become
The characteristic equation $\lambda^{2}-(\gamma+1+A)\lambda+\gamma=0$ gives the critical behavior at $|\lambda|=1$. We denote $\lambda_{w}^{\pm}:=\frac{1}{2}\left[(\gamma+1+A)\pm\sqrt{(\gamma+1+A)^{2}-4\gamma}\right]$. It can be shown that $\max_{w}|\lambda_{w}^{+}|<1$ for all parameter values under consideration. So the critical point is given by $\max_{w}|\lambda_{w}^{-}|=1$. Further, the observed period-doubling bifurcation indicates $\min_{w}\lambda_{w}^{-}=-1$, which gives $A=-2(1+\gamma)$. Since $\lambda_{w}^{-}(A)$ increases with $A$, this is achieved at $A=\min_{w}A$, i.e., $S(w):=\sum_{k=1}^{P_{c}}\cos(wk)=\csc\frac{w}{2}\sin\frac{P_{c}w}{2}\cos\frac{(P_{c}+1)w}{2}$ is minimized. In terms of the system parameters,
$\begin{split}K&=[2(1+\gamma)-K_{0}]\frac{P_{c}}{P_{c}-S_{\min}},\\ S_{\min}&:=\min_{w}\left[\csc\frac{w}{2}\sin\frac{P_{c}w}{2}\cos\frac{(P_{c}+1)w}{2}\right].\end{split}$ (2)
Note that, in the case of nearest-neighbor coupling ($P_{c}=1$), $S_{\min}=-1$ at $w=\pi$, and it gives $K_{0}=-2K+2(1+\gamma)$, which is consistent with the result in [14] (where $K\equiv 2J$).
We can infer the spatial period of the temporal period-$2$ state after the bifurcation. For example, for the patterns illustrated in Fig.1 ($N=100,P_{c}=32$), we have $\tilde{S}_{\min}=\min_{w}\left[\csc\frac{w}{2}\sin\frac{3w}{2}\cos\frac{33w}{2}\right]\approx-7.56572$ at $\tilde{w}\approx 0.13827$, by $w=\frac{2\pi l}{N}$, we have $\tilde{l}\approx 2.2$. Since the modes $w$ (and $l$) are discrete, we have denoted the continuous values by tilde. The closest integer is $l=2$ and the corresponding $S_{\min}\approx-6.94219$. The bifurcation point given by Eq.(2) is therefore $K\approx 1.31$. We refer the spatial period $l=2$ as the wavenumber.
@@
Figure 2: Snapshots of momenta for (a) $P_{c}=23$, (b) $P_{c}=17$ and (c) $P_{c}=14$. Other parameters: $\gamma=0.8$, $K_{0}=2$, $K=3$, $N=100$ and random initial conditions $(p_{j}(0),\theta_{j}(0))\in\text{Uni}[-35,35]\times\text{Uni}[0,2\pi)$, $j=1,2,...,N$. The corresponding angles behave similarly.
We note here that the wavenumber for the bifurcated state depends on the system size $N$. For a finite-size system, the eigenvalues $\lambda^{-}(l)$ are coarse-grained from the continuous ones $\lambda^{-}(\tilde{l})$, where, a coarser discretization (i.e., smaller $N$) would result in a less number of unstable eigenmodes. In a continuum limit ($N\to\infty$), the onset of instability occurs earlier than in any finite-size system. It implies that, while a smaller system realizes only one spatial pattern, therefore attracting all initial conditions in the phase space, a larger system can display multistability and realizes spatial patterns with different basin sizes. This will be discussed in the following subsection.
@@ III.2 Coherent regions and their basin sizes
[11, 21] illustrates a generic phase diagram of spatiotemporal patterns in coupled systems, where a series of decreasing-sized tongue regions highlights the coherent patterns of an increasing wavenumber and (temporal) period-doubling within each tongue. The models studied therein are coupled logistic maps, Lorenz and Rössler systems, and parameters are chosen such that the local dynamics is chaotic. Similar diagrams are also observed in other systems such as Stuart-Landau oscillators [22] and coupled Chebyshev maps (Appendix A). Here, with non-chaotic dissipative kicked rotor maps, we show that the phase diagram is qualitatively different from those in the above mentioned systems. We plot in Fig.3 the diagrams for typical wavenumbers $l$ and temporal periods $\tau$ separately.
Figure 3: Regions of (a) wavenumbers $l\leq 8$ and (b) temporal periods $\tau\leq 8$ on the parameter $(P_{c},K)$-plane, with $P_{c}\in\{1,2,...,\frac{N}{2}-1\}$ and $K\in[0.01,5]$ for $100$ equidistant values. Each pixel color is decided by the mode of $20$ trajectories. The green curve is obtained from linear stability analysis of the zero synchronized state and the green dots are the corresponding wavenumber changes. The red dots mark the instability of the temporal period-$2$ states, obtained by numerical bifurcation analysis. Other parameters: $\gamma=0.8$, $K_{0}=2$, $N=100$, and random initial conditions $(p_{j}(0),\theta_{j}(0))\in\text{Uni}[-35,35]\times\text{Uni}[0,2\pi)$, $j=1,2,...,N$. Larger-size systems show similar phase transitions, cf. Appendix B.
@@
  • Layer-I ($K\lesssim 0.4$): when the coupling is very weak, the zero homogeneous states ($l=1$, $\tau=1$) are dominant. This is well-understood from the uncoupled system that the zero fixed point has the largest basin compared to the other fixed points;
@@
  • Layer-III ($1.1\lesssim K\lesssim 1.5$): this is a transition regime where a type of temporal period-$6$ states with irregular spatial patterns dominates, but it appears only for a narrow window of $K$;
@@
Compared to the diagrams in Appendix A and in [11], first, there is no blowout bifurcation to full synchronization in the region of large coupling length $P_{c}$ and large coupling strength $K$; second, there is no tongue-shaped region with a period-doubling cascade.
@@ IV Super-long transient interface dynamics
It has been shown that for large $K_{0}$, the single rotor system has coexisting regular and chaotic attractors [14]. We now consider initially a chain of $N$ uncoupled rotors consisting of half-chain of chaotic rotors and the other half regular. It creates a domain wall between the two middle rotors $j=\frac{N}{2},\frac{N}{2}+1$, and the wall is stationary. Upon coupling, the domain wall becomes an interface, chaotic and regular motions can penetrate each other. Due to their relative basin sizes at the single-rotor level, the chaotic domain gradually encroaches the regular domain and eventually all rotors are chaotic. This has been observed in other coupled systems and has important implications in fluid dynamics [30].
@@ V Conclusion
Many interesting questions remain open. For example, we did not discuss the transitions among the Layer-I and -II, Layer-III and -IV in Fig.3(b) as they are not the main focus of this paper, but it would be interesting to study them in detail. Second, the instability of the temporal period-$2$ states – marking the onset of chaos without developing a period-doubling (or multiplying) cascade – reveals a fundamental difference from the behavior in [11] (or in Appendix A). If we compare the local map, the single rotor exhibits a period-doubling cascade to chaos [13], but the cascade occurs almost immediately and does not follow the Feigenbaum universality [14]; while the logistic map (or Chebyshev maps in Appendix A) follows the Feigenbaum universality. This may explain the difference. But the question remains: are both phase diagrams generic, and under what conditions? Moreover, while we have illustrated the relative basin sizes of the different wavenumber states, the geometry of their basins of attraction changes in the phase space remains unclear. For the super-long transient interface phenomenon, one could explore alternative initial cluster configurations, for example, setting half-chain at one fixed point and the other half at the other fixed point. Given the dominance of the chaotic basin, such configurations will eventually reach fully chaotic states, but the interface would evolve differently than presented here. Finally, we leave the analytical prediction of the sharp lifetime transition for future work.
Appendix A Coherent regions for coupled Chebyshev maps

### v2

@@ I Introduction
However, the analysis of nonlocal coupling presents significant theoretical challenges, as it eludes the simplifying symmetries of all-to-all coupling via mean-field approaches (e.g., the Kuramoto model reduces to a single order parameter), and the spatial regularity of nearest-neighbor coupling via reaction-diffusion PDEs or discrete Laplacians. Nonlocal systems, with their distance-dependent interactions – require more sophisticated approaches. The seminal work [11] demonstrated this through coherence-incoherence transitions in nonlocally coupled logistic maps, and confirmed that such a transition is universal among systems with diverse local dynamics. Inspired by this, we present an analytical and numerical study of various transitions in a nonlocally coupled system with a two-dimensional local map, specifically the dissipative kicked rotor [12, 13, 14], which has many applications in optics, engineering and quantum physics [15, 16, 17, 18]. Interestingly, despite the nonlocal coupling, linear stability analysis effectively captures the dynamics of the coupled system beyond the linearization regime; the resulting phase diagram exhibits qualitatively different structure compared to previously studied systems [11]. Analyzing bifurcations of simple states reveals transition points between coherent states of different spatial periodicity, and between coherent and incoherent states. These will be illustrated in detail.
@@ II Model
where $(p_{j}(t),\theta_{j}(t))\in(-\infty,+\infty)\times[0,2\pi)$ denote the momentum and angle of the rotor $j\in\{1,2,...N\}$ at time $t\in\mathbb{N}_{0}$, $\gamma\in(0,1)$ is the dissipation coefficient, $K_{0}>0$ is the local nonlinearity parameter, $K>0$ is the coupling strength, and $P_{c}\in\left\{1,2,..,\frac{N}{2}\right\}$ (assuming $N$ is even) is the coupling length, denoting the number of neighbors in each direction that are coupled to the rotor $j$. The momenta are set to be unbounded and non-periodic for physical relevance, though this choice does not affect the critical transitions discussed in this work.
The local map describes a dissipative kicked rotor and for a fixed $\gamma$, multiple bifurcations occur as $K_{0}$ is varied [14]: the zero state $(p,\theta)=(0,0)$ is the only attractor when $K_{0}$ is very small, and undergoes a period-doubling (PD) bifurcation at $K_{0}^{\text{PD}}=2(1+\gamma)$. Due to periodicity of $\theta$ and non-periodicity of $p$, a series of attractors is created by fold bifurcations at discrete momentum levels $p=2n\pi$, $n\in\mathbb{Z}$. These attractors further bifurcate into period-$2$ orbits and finally into chaos. The bifurcation scenarios for various values of $\gamma$ are illustrated in Fig.A1 in Appendix A, see also [14]. Smaller $\gamma$ indicates stronger damping and results in more restricted bifurcations. At $\gamma=0$, the map reduces to the one-dimensional Arnold circle map, $\theta(t+1)=\theta(t)-K_{0}\sin\theta(t)$ (mod $2\pi$), which is a fundamental model of phase locking. In contrast, it becomes the Chirikov standard map at $\gamma=1$, a classical low-dimensional example of Hamiltonian chaos where KAM tori coexist with the chaotic sea. The qualitative behavior of the local map is unaffected by $\gamma\in(0,1)$ and, as demonstrated in Appendix A, the coupled system exhibits similar critical transitions across different values of the local parameters $\gamma$ and $K_{0}$.
@@ III Spatial and Temporal Patterns
The system (1) possesses a large number of attractors due to multistability of the local map [14] and the large system size $N$. Depending on the initial conditions, the rotors self-organize into temporally periodic patterns with or without a distinguishable spatial period. Consequently, for any given values of the coupling parameters $(P_{c},K)$, multiple attractors can coexist. However, we notice that for a specific parameter region, the dominant attractors can be characterized by low temporal and spatial periods.
In this section, we focus on the following three issues. The first one is the transition from the stationary, fully synchronized state to a coherent state of temporal period-$2$, which can be explained analytically through linear stability analysis. A wavenumber will be defined to characterize the spatial periodicity. The second issue is to study coherent regions on the parameter $(P_{c},K)$-plane, particularly the regions with low spatial and temporal periods. The transition curves can be obtained from stability analysis of the temporal period-$2$ states. The third issue concerns the coexistence of the temporal period-$2$ states with different wavenumbers, and we numerically examine how their relative basin sizes vary with the coupling length $P_{c}$.
Throughout this section, purely random initial conditions are used for numerical simulations: $(p_{j}(0),\theta_{j}(0))\in\text{Uni}[-p_{0},p_{0}]\times\text{Uni}[0,2\pi)$, where $p_{0}>2\pi$ is constant (the specific value of $p_{0}$ does not affect the qualitative results).
@@ III.1 Transition from zero homogeneous states
The full phase space of the system (1) is a product of $N$ infinite cylinders $(p_{j},\theta_{j})\in(-\infty,+\infty)\times[0,2\pi)$, $\forall j=1,2,...,N$. For $K_{0}$ and $K$ both small, trajectories initially close to $\boldsymbol{p}:=(p_{1},p_{2},...,p_{N})=\boldsymbol{0}$ will converge to the zero synchronized state $(\boldsymbol{p},\boldsymbol{\theta})=(\boldsymbol{0},\boldsymbol{0})$ and it appears as the only attractor. However, due to multistability of the local map, more attractors can be observed when we consider longer truncated cylinders (i.e., $\boldsymbol{p}$ further away from $\boldsymbol{0}$), and the notion of the zero synchronized state is extended to a state in which most rotors are homogeneous but with a few exceptions; these exceptional rotors locate randomly depending on the initial condition, and stay at the nonzero fixed points of the local map. We refer such a state a zero homogeneous state. An example is shown in Fig.1(a): for $K_{0}=2$ the local map has three fixed points $p=0,\pm 2\pi$, where $p=0$ has the largest basin and $p=\pm 2\pi$ have equal basin sizes [14].
Figure 1: Snapshots of momenta for (a) $K=1.3$, (b) $K=1.4$ and (c) $K=3$. Other parameters: $\gamma=0.8$, $K_{0}=2$, $P_{c}=32$, $N=100$ and random initial conditions $(p_{j}(0),\theta_{j}(0))\in\text{Uni}[-35,35]\times\text{Uni}[0,2\pi)$, j=1,2,,Nj=1,2,...,N. The corresponding angles behave similarly.
@@
This period-doubling bifurcation can be understood analytically via linear stability analysis, giving the critical parameter values at which the Jacobian eigenvalues first cross the unit circle (details see Appendix B):
$\begin{split}K&=[2(1+\gamma)-K_{0}]\frac{P_{c}}{P_{c}-S_{\min}},\\ S_{\min}&:=\min_{w}\left[\csc\frac{w}{2}\sin\frac{P_{c}w}{2}\cos\frac{(P_{c}+1)w}{2}\right],\end{split}$ (2)
where $w=\frac{2\pi l}{N}$, l=0,1,,N1l=0,1,...,N-1 (for periodic boundary conditions). Note that, in the case of nearest-neighbor coupling ($P_{c}=1$), $S_{\min}=-1$ at $w=\pi$, giving the critical parameter curve $K_{0}=-2K+2(1+\gamma)$, which is consistent with the result in [14] (where $K\equiv 2J$); moreover, $w=\pi$ being the most unstable mode indicates that every two consecutive rotors form a wave packet, referred to as the alternating state in [14].
We can also infer the spatial period of the bifurcated state by the value of $w$ (thus $l$) at criticality. For example, for the patterns illustrated in Fig.1 ($N=100,P_{c}=32$), we have $\tilde{S}_{\min}=\min_{w}\left[\csc\frac{w}{2}\sin\frac{3w}{2}\cos\frac{33w}{2}\right]\approx-7.56572$ at $\tilde{w}\approx 0.13827$, by $w=\frac{2\pi l}{N}$, we have $\tilde{l}\approx 2.2$. Since the modes $w$ (and $l$) are discrete, we have denoted the continuous values by tilde. The closest integer is $l=2$ and the corresponding $S_{\min}\approx-6.94219$. The bifurcation point given by Eq.(2) is therefore $K\approx 1.31$. We refer the spatial period $l=2$ as the wavenumber.
@@
Figure 2: Snapshots of momenta for (a) $P_{c}=23$, (b) $P_{c}=17$ and (c) $P_{c}=14$. Other parameters: $\gamma=0.8$, $K_{0}=2$, $K=3$, $N=100$ and random initial conditions $(p_{j}(0),\theta_{j}(0))\in\text{Uni}[-35,35]\times\text{Uni}[0,2\pi)$, j=1,2,,Nj=1,2,...,N. The corresponding angles behave similarly.
[…]
@@ III.2 Coherent regions and their basin sizes
[11, 25] illustrates a generic phase diagram of spatiotemporal patterns in coupled systems, where a series of decreasing-sized tongue regions highlights the coherent patterns of an increasing wavenumber and (temporal) period-doubling within each tongue. The models studied therein are coupled logistic maps, Lorenz and Rössler systems, and parameters are chosen such that the local dynamics is chaotic. Similar diagrams are also observed in other systems such as Stuart-Landau oscillators [26] and coupled Chebyshev maps (Appendix C). Here, with non-chaotic dissipative kicked rotor maps, we present that the phase diagram is qualitatively different from those in the above mentioned systems. Specifically, (i) there is no blowout bifurcation to full synchronization in the regime of large coupling length $P_{c}$ and large coupling strength $K$, (ii) there is no tongue-shaped region exhibiting a period-doubling cascade, and (iii) the coherence-incoherence transition does not appear to transit through partially synchronized (or chimera) states, or at least they are not dominant for any parameter values.
We plot the diagrams in Fig.3 for typical wavenumbers $l$ and temporal periods $\tau$ separately, where $l$ is determined by a numerical Fourier transform and $\tau$ by periodicity of the time series (up to numerical precision).
Figure 3: Regions of typical (a) wavenumbers $l\leq 8$ and (b) temporal periods $\tau\leq 8$ on the parameter $(P_{c},K)$-plane, with $P_{c}\in\{1,2,...,\frac{N}{2}-1\}$ and $K\in[0.01,5]$ for $100$ equidistant values. Each pixel color is decided by the mode over 2020 trajectories. The green curve is obtained from linear stability analysis of the zero synchronized state and the green dots are the corresponding wavenumber changes. The red dots mark the instability of the temporal period-$2$ states, obtained by numerical bifurcation analysis. Other parameters: $\gamma=0.8$, $K_{0}=2$, $N=100$, and random initial conditions $(p_{j}(0),\theta_{j}(0))\in\text{Uni}[-35,35]\times\text{Uni}[0,2\pi)$, j=1,2,,Nj=1,2,...,N. Larger-size systems show similar phase transitions, cf. Appendix D.
@@
  • Layer-I ($K\lesssim 0.4$): when the coupling is very weak, the zero homogeneous states ($l=0$, $\tau=1$) are dominant. This is well-understood from the uncoupled system that the zero fixed point has the largest basin compared to the other fixed points;
@@
  • Layer-III ($1.1\lesssim K\lesssim 1.5$): this is a transition regime where a type of temporal period-$6$ states with irregular spatial patterns dominates, but it appears only for a narrow window of $K$; note that here many patterns coexist, the example of $\tau=2$ shown in Fig.1(b) are stable but its basin is relatively small compared to the $\tau=6$ states;
@@ IV Super-long transient interface dynamics
It has been shown that for large $K_{0}$, the single-rotor system has coexistence of chaos and tiny regular regions [14], which is also a common feature in Hamiltonian systems such as Lorentz gases [34]. We now consider initially a chain of $N$ uncoupled rotors consisting of half-chain of chaotic rotors and the other half regular. It creates a domain wall between the two middle rotors $j=\frac{N}{2},\frac{N}{2}+1$, and the wall is stationary. Upon coupling, the domain wall becomes an interface, chaotic and regular motions can penetrate each other. Due to their relative basin sizes at the single-rotor level, the chaotic domain gradually encroaches the regular domain and eventually all rotors are chaotic. This has been observed in other coupled systems and has important implications in fluid dynamics [35].
@@ V Conclusion
Many interesting questions remain open. For example, we did not discuss the transitions among the Layer-I and -II, Layer-III and -IV in Fig.3(b) as they are not the main focus of this paper, but it would be interesting to study them in detail. Second, the instability of the temporal period-$2$ states – marking the onset of chaos without developing a period-doubling (or multiplying) cascade – reveals a fundamental difference from the behavior in [11] (or in Appendix C). If we compare the local map, the single rotor exhibits a period-doubling cascade to chaos [13], but the cascade occurs almost immediately and does not follow the Feigenbaum universality [14]; while the logistic map (or Chebyshev maps in Appendix C) follows the Feigenbaum universality. This may explain the difference. But the question remains: are both phase diagrams generic, and under what conditions? Moreover, while we have illustrated the relative basin sizes of the different wavenumber states, the geometry of their basins of attraction changes in the phase space remains unclear. For the super-long transient interface phenomenon, one could explore alternative initial cluster configurations, for example, setting half-chain at one fixed point and the other half at the other fixed point. Given the dominance of the chaotic basin, such configurations will eventually reach fully chaotic states, but the interface would evolve differently than presented here. Finally, we leave the analytical prediction of the sharp lifetime transition for future work.
Acknowledgments The author would like to thanks the two anonymous referees for their suggestions, which helped improve the quality and clarity of this manuscript.
Appendix A Influence of the dissipation coefficient $\gamma$
We illustrate here that the qualitative dynamics of the single-rotor model and the existence of the coherent region in the coupled system are not influenced by the value of the dissipation coefficient $\gamma\in(0,1)$.
In Fig.A1, we plot, for three different values of $\gamma$, the bifurcations in the single-rotor model for the momentum $p$ (upper panel) and the angle $\theta$ (lower panel). Stronger dissipation (i.e., smaller $\gamma$) reduces the accessible momentum range and limits the number of possible bifurcating branches, but it does not alter the qualitative branching structure of period-doubling and fold bifurcations [14].
Figure A1: Bifurcations in the single dissipative rotor with (a)-(b) $\gamma=0.55$, (c)-(d) $\gamma=0.8$ and (e)-(f) $\gamma=0.95$. The upper panel shows for momentum $p$ and the lower panel for angle $\theta$.
Figure A2: Regions of typical wavenumbers $l\leq 8$ and temporal periods $\tau\leq 8$ on the parameter $(P_{c},K)$-plane. Upper panel: $(\gamma,K_{0})=(0.55,2)$ and lower panel: $(\gamma,K_{0})=(0.95,2)$. Other numerical settings are as in Fig.3.
On the other hand, the dynamics of the coupled system shown in Fig.3 is generic for various values of $\gamma$. Two examples are illustrated in Fig.A2: compared to $\gamma=0.8$ in the main text, the coherence region (Layer-IV in Fig.3) is present for both stronger ($\gamma=0.55$) and weaker ($\gamma=0.95$) dissipation, which is characterized by temporal period $\tau=2$ and an increasing wavenumber $l$ over progressively narrower sub-intervals of the coupling length $P_{c}$. The only reason we use $\gamma=0.8$ together with $K_{0}=2$ in Sec.III is that, as shown in Figs.A1(c)-(d), the single rotor exhibits a relatively simple attractor structure, namely the three fixed points at $p=0,\pm 2\pi$. This simplicity carries over to the weakly coupled regime, allowing a clear demonstration of the critical transitions.
Appendix B Linear stability and behavior of the Jacobian eigenvalues
Under the assumption of slow angle variations, the equations of momenta in Eq.(1) can be linearized as
By a Fourier transform $p_{j}(t)=\sum_{w}P_{w}(t)e^{iwj}$, $\theta_{j}(t)=\sum_{w}\Theta_{w}(t)e^{iwj}$, where $w=\frac{2\pi l}{N}$ and l=0,1,,N1l=0,1,...,N-1 (for periodic boundary conditions) we have, for each Fourier mode $w$,
The linearized equations of motion thus become
The characteristic equation $\lambda^{2}-(\gamma+1+A)\lambda+\gamma=0$ gives the critical behavior at $|\lambda|=1$. We denote
$\lambda_{w}^{\pm}:=\frac{1}{2}\left[(\gamma+1+A)\pm\sqrt{(\gamma+1+A)^{2}-4\gamma}\right].$
The maximum of $|\lambda_{w}^{+}|$ occurs when $A=\max_{w}A=-K_{0}$ at $w=0$, and we have $\max_{w}|\lambda_{w}^{+}|=\frac{1}{2}[(\gamma+1-K_{0})+\sqrt{(\gamma+1-K_{0})^{2}-4\gamma}]<1+(\gamma-K_{0})<1$ for all parameter values under consideration ($\gamma\in(0,1)$ and $K_{0}>1$). So the critical point is given by $\max_{w}|\lambda_{w}^{-}|=1$. Further, the observed period-doubling bifurcation indicates $\min_{w}\lambda_{w}^{-}=-1$, which gives $A=-2(1+\gamma)$. Since $\lambda_{w}^{-}(A)$ increases with $A$, this is achieved at $A=\min_{w}A$, i.e., $S(w):=\sum_{k=1}^{P_{c}}\cos(wk)=\csc\frac{w}{2}\sin\frac{P_{c}w}{2}\cos\frac{(P_{c}+1)w}{2}$ is minimized. Let us denote $S_{\min}:=\min_{w}\left[\csc\frac{w}{2}\sin\frac{P_{c}w}{2}\cos\frac{(P_{c}+1)w}{2}\right]$, then the criticality occurs at
$2(1+\gamma)=-(K_{0}+K)+\frac{K}{P_{c}}S_{\min},$
or equivalently,
$K=[2(1+\gamma)-K_{0}]\frac{P_{c}}{P_{c}-S_{\min}}.$
For the examples showed in Fig.2, we plot the eigenvalues $\lambda_{w}^{-}$ in Fig.A3 before ($K=1.3$) and after ($K=1.4$) the period-doubling bifurcation. Fig.A3(c) illustrates that $l=2$ is the most unstable discrete eigenmode, corresponding to a spatial pattern of wavenumber $l=2$.
Figure A3: Eigenvalues $\lambda^{-}$ before (a) $K=1.3$ and after (b)-(c) $K=1.4$ the period-doubling bifurcation; (c) shows $|\lambda^{-}|$ in continuous wavenumbers $\tilde{l}$ (blue) and in discrete $l$ (orange). Other parameter values: $\gamma=0.8$, $K_{0}=2$, $P_{c}=32$ and $N=100$.
Beyond the first instability, additional unstable modes emerge in the eigenspectrum. Fig.A4 shows the cases after two and three successive crossings of the unit circle.
Figure A4: $|\lambda^{-}|$ in continuous wavenumbers $\tilde{l}$ (blue) and in discrete $l$ (orange): (a) $K=1.48$ and (b) $K=1.52$. Other parameter values: $\gamma=0.8$, $K_{0}=2$, $P_{c}=32$ and $N=100$.
Fig.A5 presents the influence of the coupling length $P_{c}$ on the unstable modes: for a fixed $K$, decreasing $P_{c}$ leads to larger unstable wavenumbers $l$. In particular, the nearest-neighbor coupling ($P_{c}=1$) yields $l=\frac{N}{2}$, corresponding to a spatially alternating pattern.
Figure A5: $|\lambda^{-}|$ in continuous wavenumbers $\tilde{l}$ for different coupling lengths $P_{c}$. Other parameter values: $\gamma=0.8$, $K_{0}=2$, $K=1.4$ and $N=100$.
Appendix C Coherent regions for coupled Chebyshev maps